# Friederike Becker1, Fabian Krüger1 and Melanie Schienle1,2

### v2

@@ 1 Introduction
Macroeconomic forecasts feature prominently in economic policy debates, and outlooks are often compared in the cross-section of economically similar countries across the globe. While the academic literature on macroeconomic forecasting has long appreciated the importance of measuring and communicating forecast uncertainty (as represented, e.g., by distributions and prediction intervals), the broader policy debate is still dominated by point forecasts. This may partly be due to the fact that for many countries including the Group of 7 economies (henceforth G7, including Canada, France, Germany, Italy, Japan, the United Kingdom and the United States) real-time forecast distributions are not readily available. In case of the G7 and recent editions of macroeconomic forecasting reports issued by their central banks, four of the seven reports did not explicitly quantify uncertainty. Table 1 and the appendix provide details. Note that among the many economic organizations that issue forecasts, central banks take a leading role in terms of both societal impact and technical sophistication. If anything, we thus expect their coverage of forecast uncertainty to be more advanced than that of other economic organizations. Hence, quantitative information on macroeconomic forecast uncertainty is often unavailable, even for well-informed readers who invest the time to consult rather specialized documents.
@@
Italy Banca d’Italia no
@@
Table 1: State of uncertainty quantification for macroeconomic forecasts that are issued by the central banks of the G7 countries. For each central bank, we consulted the “flagship” report containing forecasts for inflation and real GDP growth, often titled “Macroeconomic Projections”, “Economic Outlook” or “Monetary Policy Report” (depending on availability) and checked whether that document’s centrally communicated forecasts are accompanied by an explicit quantification of uncertainty. For more details on the reports and our protocol for extracting the information contained therein, please refer to the appendix.
Based on work by Adams et al. (2021), a recent project by the Federal Reserve Bank of New York (2024) provides probabilistic forecasts of real GDP growth, inflation and unemployment in the US. While similar in spirit, our forecast distributions are based on public data (rather than proprietary survey forecast data), and cover all G7 economies. In line with the IMF’s World Economic Outlook, our forecasts are bi-annual (rather than monthly). Furthermore, our methodology is somewhat simpler, in that we do not attempt to predict the distribution of forecast errors by means of additional variables. Schick (2024) incorporates SPF forecasts into an AR-GARCH model to forecast quantiles of U.S. GDP growth. Reifschneider and Tulip (2019) describe prediction intervals presented by the U.S. Federal Reserve, based on root mean squared forecasting errors. The intervals’ coverage level depends on the distribution of forecast errors, amounting to roughly 70 percent under a normal distribution. Elder et al. (2005) evaluate prediction intervals implied by the Bank of England’s ‘fan charts’, which are based on a skewed distribution reflecting the Bank’s judgment. Compared to these parametric approaches, our proposed technique considers empirical quantiles of absolute forecast errors, thus avoiding restrictive functional form assumptions. The robustness of quantiles is an important advantage in turbulent times like the recent corona pandemic, whose extreme observations are handled plausibly and automatically by quantiles. By contrast, non-robust estimation methods based on squared errors are dominated by these observations. This leads to difficult questions about how to handle extreme observations in practice, with many studies resorting to ad-hoc choices of the sample period. See Lenza and Primiceri (2022) and Knüppel et al. (2023, Section 5) for further discussion.
@@
Finally, our paper relates to a growing body of academic literature on how to construct forecast distributions of macroeconomic variables. Our proposed approach intends to be as simple as possible (in terms of data requirements and statistical techniques), subject to being reasonably competitive in terms of statistical performance. Studies like Clements (2010), Krüger (2017), Ganics et al. (2022) and Krüger and Plett (2024) indicate that forecast distributions based on past point forecast errors perform similar to, or better than, subjective histogram-type forecasts as provided by the SPF and ECB-SPF. These findings motivate our use of the IMF WEO as an external point forecast, together with a suitable set of historical forecast errors. See Qu et al. (2024) for recent evidence on the good performance of IMF point forecasts relative to private sector survey forecasts. Compared to studies that create ‘self-contained’ forecast distributions based on multivariate time series data (see e.g. Clark et al., 2024, and the references therein), our use of the IMF as an external point forecast greatly reduces data requirements and modeling complexity. For our purposes, this simplification outweighs the inherent benefits of producing a self-contained forecast. More broadly, the principle of constructing forecast distributions based on a history of point forecasts and associated realizations has proven successful in many empirical contexts across the disciplines. Similar to our main approach, Kath and Ziel (2021) include a model that constructs forecast intervals based on quantiles of absolute point forecast errors, in the context of high-frequency electricity spot price data. Walz et al. (2024) and Angelopoulos et al. (2023) provide further discussion in the context of meteorology and machine learning, respectively.
@@
The remainder of this paper is organized as follows. Section 2 contains methodology for computing empirical quantiles in the current setup. In particular, we discuss techniques for ensuring that forecasts are coherent across multiple horizons, as well as assumptions about the (a)symmetry of forecast errors. Section 3 describes relevant benchmark methods and presents empirical results on the forecasting performance of the proposed method. In our empirical analysis, we show that the method yields calibrated prediction intervals (with coverage close to its nominal level), and performs at least as well as the benchmarks in terms of the interval score, a statistical loss function for prediction intervals. Section 4 concludes. The appendix contains details on Table 1, as well as further empirical results and robustness checks. Replication materials for the paper are available through Becker et al. (2023).
@@ 2.1 Constructing Prediction Intervals
with $c^{u,\tau}>c^{l,\tau}$. Note that $c^{l,\tau}$ can but does not have to be negative. To get a finer representation of the predictive distribution, one can construct these intervals for multiple levels of $\tau$. Given $\tau$, we obtain values for the upper and lower prediction bands $c^{u,\tau}\text{ and }c^{l,\tau}$ from past data on forecasts and realizations, according to the following methodology.
@@ 2.2 Absolute versus Raw Errors
Table 2 summarizes our considerations for deciding between absolute versus directional errors. In view of these arguments, we opt to use absolute errors as our default option. Of course, this choice reflects a weighting of the different criteria mentioned in Table 2, which is necessarily subjective.
@@ 2.3 Assessing Forecast Accuracy
In assessing the quality of our forecasts, we follow the paradigm to ‘maximize sharpness subject to calibration’ postulated by Gneiting et al. (2007). Sharpness means that the forecast distribution should be as narrow as possible. Calibration means that the forecast distribution should be coherent with observed outcomes. In practice, there is a clear trade-off between both objectives. For example, a forecast distribution with all probability mass in one point (indicating no uncertainty) would be perfect in terms of sharpness, but likely poor in terms of calibration. In order to assess the calibration of probabilistic forecasts in an interval format, a simple method is to compare the empirical and nominal interval coverage rates. That is, a $\tau\%$ coverage interval should over time cover roughly $\tau\%$ of observations. As argued by Raftery (2016), forecasts users tend to view calibration as a crucial feature of a trustworthy forecast. A proper scoring rule can then be used to evaluate the forecast sharpness in conjunction with its calibration properties. Briefly, a scoring rule is called proper if it encourages forecasters to state what they think is the correct prediction (Winkler, 1996; Gneiting and Raftery, 2007). Given the format of our forecast, the interval score (Gneiting and Raftery, 2007) is a natural choice of scoring rule:
@@
for a given confidence level $\tau$ and the corresponding prediction interval $[l,u]$ implied by the forecast distribution $F$. Here we consider scores in negative orientation, i.e., smaller scores correspond to better forecasts. The first summand of Equation (7) can be interpreted as a penalty for dispersion (i.e., lack of sharpness) of the forecast distribution, corresponding to wide prediction intervals, such that $u-l$ is large. The second summand can be seen as a penalty for overprediction, such that the outcome $y$ is smaller than the interval’s lower endpoint $l$. Conversely, the third summand of (7) represents a penalty for underprediction.
Given that (weighted) sums of proper scoring rules are again proper (Gneiting and Raftery, 2007), we can average scores for different levels of $\tau=\tau_{1},\ldots,\tau_{K}$ in order to obtain a summary measure of forecast performance. In the following, we consider a weighted average with weights $w_{1},...,w_{K}$ as proposed by Bracher et al. (2021a). Specifically, we set $K=2$, $\tau_{1}=0.5$ and $\tau_{2}=0.8$, and thereby $w_{1}=\frac{1-\tau_{1}}{2}=0.25$ and $w_{2}=\frac{1-\tau_{2}}{2}=0.1$. Since we focus on generating prediction intervals, we do not incorporate the point (median) prediction into the interval score. The latter is externally provided by the IMF in our empirical setup.
@@ 3.1 Data and Forecast Targets
The WEO’s merit relative to other forecast sources has been established in studies such as Timmermann (2007) and Qu et al. (2024), which is why we consider it a promising candidate for this analysis. Started in 1990, the WEO has issued bi-annual forecasts for multiple (up to six) years into the future, for several variables including real GDP growth and inflation (International Monetary Fund, 2024). The code that downloads and processes the current WEO dataset from the IMF website is available in a public GitHub repository (Becker et al., 2023).
@@
We focus on 80% and 50% prediction intervals. As argued by Raftery (2016), an 80% interval satisfies the intuitive notion that a prediction interval should ‘typically’ cover the realizing outcome, while often being reasonably short (i.e., sharp) in practice. The $50\%$ interval is additionally constructed to provide further characterization of the forecast distribution for interested users, and to serve as a further check of the validity of the methodology.
@@ 3.2 Choice of Tuning Parameters
Table 4 and Figure 6 in the appendix summarize the training sample performance of various combinations of the tuning parameters. Based on these results, and the considerations in Section 2, we use a rolling window of $R=11$ observations, the absolute error method, and the default empirical quantile method (type = 7) [^1] Note that for $R=11$ and the absolute error method, this is also equivalent to the inverse of the empirical distribution function, that is, type = 1. As we do not estimate quantiles at the tails of the distribution the arguments of Hyndman and Fan (1996) in favor of the type = 8 method do not apply for our setting. of the quantile function in R. Our training sample results indicate that shorter window lengths $R<11$ perform worse in terms of the IS, see Figure 6 in the appendix. The difference in performance is particularly pronounced for very short window lengths $74\leq R<$ 7 and at longer forecast horizons. By contrast, the benefit of increasing $R$ from $8$ to $11$ is remarkably small in all empirical setups (absolute and directional errors, all forecast horizons). Hence the marginal benefit of increasing the sample size seems to vanish rapidly as $R$ increases. We exclude choices $R>11$ since we cannot evaluate their performance given the time span of our training data, while an extension of the training data would have reduced the time span of our holdout sample, thus limiting our ability to assess out-of-sample performance. From a statistical perspective, using a short window length $R$ ensures adaptability to potential structural breaks (c.f. Inoue et al., 2017). We find that coverage rates on the training data are generally more favorable for the absolute than for the directional error method, sometimes on the order of about 10 percentage points (see Table 4 in the appendix), while interval scores are similar.
@@
Figure 2: Illustration of the process that we use to construct prediction intervals at the $50\%$ and $80\%$ confidence levels, using point forecasts and realized values from the International Monetary Fund World Economic Outlook forecast data.
@@ 3.3 Benchmark Models
In order to put the performance of our proposed method in perspective, we compare it to three benchmarks. For the first benchmark (denoted AR), we compute point predictions from an autoregressive model of order one, estimated via ordinary least squares. We fit the model to quarterly data on the predictand (inflation or GDP growth) obtained from the OECD, and construct a forecast of the annual series as a weighted sum of quarterly forecasts. See Section B.1 in the appendix for details. We then construct prediction intervals from empirical quantiles of the model’s past forecast errors, analogous to our proposed procedure for the IMF forecasts.
For the second benchmark (denoted BVAR), we consider a Bayesian vector autoregressive model specification proposed by Primiceri (2005) and Del Negro and Primiceri (2015), and implemented in the R package bvarsv (Krüger, 2015). In our main results, we consider two quarterly system variables (inflation and GDP growth) obtained from the OECD. In a robustness check described below, we add an index of financial stress as a third variable. Section B.1 in the appendix provides details. As for the IMF forecasts (and for the AR model), we then construct prediction intervals from empirical quantiles of the BVAR’s past forecast errors.
@@
For the third benchmark (denoted BVAR-direct), we use the same model specification as BVAR, but compute prediction intervals directly from the model’s forecast distribution. That is, the forecast distribution is based on the model’s parametric assumptions, rather than on the empirical distribution of the model’s past point forecast errors.
In order to roughly match the timing of IMF forecasts, we equip the benchmarks for the spring forecasts with data up until the first quarter, and the benchmarks for the fall forecasts with data up until the third quarter.
The AR and BVAR benchmarks are analogous to our proposed methodology (based on IMF forecasts), except that they use a different point forecast. Hence they allow to assess the competitiveness of the IMF’s point forecasts in our setup. While AR is simple and transparent, BVAR is much more flexible, allowing for time-varying parameters. The latter comes at the cost of increased complexity (in particular, choice of prior parameters and simulation of the posterior distribution based estimation via Markov chain Monte Carlo methods). In contrast to AR and BVAR, BVAR-direct is fully based on parametric modeling assumptions.

### v3

@@ 1 Introduction
Macroeconomic forecasts feature prominently in economic policy debates, and outlooks are often compared in the cross-section of economically similar countries across the globe. While the academic literature on macroeconomic forecasting has long appreciated the importance of measuring and communicating forecast uncertainty (as represented, e.g., by distributions and prediction intervals), the broader policy debate is still dominated by point forecasts. This may partly be due to the fact that for many countries including the Group of 7 economies (henceforth G7, including Canada, France, Germany, Italy, Japan, the United Kingdom and the United States), real-time forecast distributions are not readily available. In case of the G7 and recent editions of macroeconomic forecasting reports issued by their central banks, three of the seven reports did not explicitly quantify uncertainty. Table 1 and Section A of the supplement provide details. Among the many economic organizations that issue forecasts, central banks take a leading role in terms of both societal impact and technical sophistication. If anything, we thus expect their coverage of forecast uncertainty to be more advanced than that of other economic organizations. Hence, quantitative information on macroeconomic forecast uncertainty is often unavailable, even for well-informed readers who invest the time to consult rather specialized documents.
@@
Italy Banca d’Italia yes
@@
Table 1: State of uncertainty quantification for macroeconomic forecasts that are issued by the central banks of the G7 countries. For each central bank, we consulted the “flagship” report containing forecasts for inflation and real GDP growth, often titled “Macroeconomic Projections”, “Economic Outlook” or “Monetary Policy Report” (depending on availability) and checked whether that document’s centrally communicated forecasts are accompanied by an explicit quantification of uncertainty. For more details on the reports and our protocol for extracting the information contained therein, see Section A in the supplement.
Based on work by Adams et al. (2021), a recent project by the Federal Reserve Bank of New York (2024) provides probabilistic forecasts of real GDP growth, inflation and unemployment in the U.S. While similar in spirit, our forecast distributions are based on public data (rather than proprietary survey forecast data), and cover all G7 economies. In line with the IMF’s World Economic Outlook, our forecasts are bi-annual (rather than monthly). Furthermore, our methodology is somewhat simpler, in that we do not attempt to predict the distribution of forecast errors by means of additional variables. Schick (2024) incorporates SPF forecasts into an AR-GARCH model to forecast quantiles of U.S. GDP growth. Reifschneider and Tulip (2019) describe prediction intervals presented by the U.S. Federal Reserve, based on root mean squared forecasting errors. The intervals’ coverage level depends on the distribution of forecast errors, amounting to roughly 70 percent under a normal distribution. Elder et al. (2005) evaluate prediction intervals implied by the Bank of England’s ‘fan charts’, which are based on a skewed distribution reflecting the Bank’s judgment. Compared to these parametric approaches, our proposed technique considers empirical quantiles of absolute forecast errors, thus avoiding restrictive functional form assumptions. The robustness of quantiles is an important advantage in turbulent times like the Covid-19 pandemic, whose extreme observations are handled plausibly and automatically by quantiles. By contrast, non-robust estimation methods based on squared errors are dominated by these observations. This leads to difficult questions about how to handle extreme observations in practice, with many studies resorting to ad-hoc choices of the sample period. See Lenza and Primiceri (2022) and Knüppel et al. (2023, Section 5) for further discussion.
@@
Finally, our paper relates to a growing body of academic literature on how to construct forecast distributions of macroeconomic variables. Our proposed approach intends to be as simple as possible (in terms of data requirements and statistical techniques), subject to being reasonably competitive in terms of statistical performance. Studies like Clements (2010), Krüger (2017), Ganics et al. (2022) and Krüger and Plett (2024) indicate that forecast distributions based on past point forecast errors perform similar to, or better than, subjective histogram-type forecasts as provided by the SPF and ECB-SPF. These findings motivate our use of the IMF WEO as an external point forecast, together with a suitable set of historical forecast errors. See Qu et al. (2024) for recent evidence on the good performance of IMF point forecasts relative to private sector survey forecasts. The IMF point forecasts are survey based and may thus incorporate judgmental information that is hard to incorporate into formal time series models (e.g., information on fiscal or monetary policy announcements, or recent releases of weekly or monthly macroeconomic time series). For U.S. data, studies such as Faust and Wright (2009) and Krüger et al. (2017) have found that judgmental information is most helpful at short forecast horizons (where survey forecasts of GDP growth and inflation are hard to beat), while being less effective at longer horizons. By using the IMF point forecast, our method exploits the potential benefits of judgmental information in terms of point forecasting. At the same time, and as motivated above, our assessment of forecast uncertainty is purely statistical. Compared to studies that create ‘self-contained’ forecast distributions based on multivariate time series data (see e.g. Clark et al., 2024, and the references therein), our use of the IMF WEO as an external point forecast greatly reduces data requirements and modeling complexity. For our purposes, this simplification outweighs the inherent benefits of producing a self-contained forecast. More broadly, the principle of constructing forecast distributions based on a history of point forecasts and associated realizations has proven successful in many empirical contexts across the disciplines. Similar to our main approach, Kath and Ziel (2021) include a model that constructs forecast intervals based on quantiles of absolute point forecast errors, in the context of high-frequency electricity spot price data. Walz et al. (2024) and Angelopoulos et al. (2023) provide further discussion in the context of meteorology and machine learning, respectively. While we focus on forecast errors constructed from IMF point forecasts, the methodological questions we discuss (in particular, the use of absolute versus raw forecast errors, and the concrete specification of window method for selecting the relevant training data) arise similarly when using any other source of point forecasts. These other sources include statistical or machine learning methods for which point forecasts are often more easily accessible via software packages than quantile or distribution forecasts. This paper hence offers a framework for designing error-based forecast distributions in empirical macroeconomics.
@@
The remainder of this paper is organized as follows. Section 2 contains methodology for computing empirical quantiles in the current setup. In particular, we discuss techniques for ensuring that forecasts are coherent across multiple horizons, as well as assumptions about the (a)symmetry of forecast errors. Section 3 describes relevant benchmark methods and presents empirical results on the forecasting performance of the proposed method. In our empirical analysis, we show that the method yields calibrated prediction intervals (with coverage close to its nominal level), and overall performs similarly to or better than the benchmarks in terms of the interval score, a statistical loss function for prediction intervals. Section 4 concludes. The supplement contains details on Table 1, as well as further empirical results and robustness checks. Replication materials for the paper are available through Becker et al. (2023).
@@ 2.1 Constructing Prediction Intervals
with $c^{u,\tau}>c^{l,\tau}$. Note that $c^{l,\tau}$ can, but does not have to, be negative. To get a finer representation of the predictive distribution, one can construct these intervals for multiple levels of $\tau$. Given $\tau$, we obtain values for the upper and lower prediction bands $c^{u,\tau}\text{ and }c^{l,\tau}$ from past data on forecasts and realizations, according to the following methodology.
@@ 2.2 Absolute versus Raw Errors
Table 2 summarizes our general considerations for deciding between absolute versus directional errors. Section B.3.1 in the supplement also presents illustrative evidence of empirical advantages of absolute versus directional errors for our specific small sample setup. In view of these arguments, we opt to use absolute errors as our default option. Of course, this choice reflects a weighting of the different criteria mentioned in Table 2, which is necessarily subjective.
@@ 2.3 Assessing Forecast Accuracy
In assessing the quality of our forecasts, we follow the paradigm to ‘maximize sharpness subject to calibration’ postulated by Gneiting et al. (2007). Sharpness means that the forecast distribution should be as narrow as possible. Calibration means that the forecast distribution should be coherent with observed outcomes. In practice, there is a clear trade-off between both objectives. For example, a forecast distribution with all probability mass in one point (indicating no uncertainty) would be perfect in terms of sharpness, but likely poor in terms of calibration. In order to assess the calibration of probabilistic forecasts in an interval format, a simple method is to compare the empirical and nominal interval coverage rates. That is, a $\tau\%$ coverage interval should over time cover roughly $\tau\%$ of observations. As argued by Raftery (2016), forecasts users tend to view calibration as a crucial feature of a trustworthy forecast in practice. A proper scoring rule can then be used to evaluate the forecast sharpness in conjunction with its calibration properties. Briefly, a scoring rule is called proper if it encourages forecasters to state what they think is the correct prediction (Winkler, 1996; Gneiting and Raftery, 2007). Given the format of our forecast, the interval score (Gneiting and Raftery, 2007) is a natural choice of scoring rule:
@@
for a given confidence level $\tau$ and the corresponding prediction interval $[l,u]$ implied by the forecast distribution $F$. Here we consider scores in negative orientation, i.e., smaller scores correspond to better forecasts. The first summand of Equation (7) can be interpreted as a penalty for dispersion (i.e., lack of sharpness) of the forecast distribution, corresponding to wide prediction intervals, such that $u-l$ is large. The second summand can be seen as a penalty for overprediction, such that the outcome $y$ is smaller than the interval’s lower endpoint $l$. Conversely, the third summand of (7) represents a penalty for underprediction. Given that (weighted) sums of proper scoring rules are again proper (Gneiting and Raftery, 2007), we can add scores for different levels of $\tau=\tau_{1},\ldots,\tau_{K}$ in order to obtain a summary measure of forecast performance. In the following, we consider a weighted sum with weights $w_{1},...,w_{K}$ as proposed by Bracher et al. (2021a). Specifically, we set $K=2$, $\tau_{1}=0.5$ and $\tau_{2}=0.8$, and thereby $w_{1}=\frac{1-\tau_{1}}{2}=0.25$ and $w_{2}=\frac{1-\tau_{2}}{2}=0.1$. Since we focus on generating prediction intervals, we do not incorporate the point (median) prediction into the interval score. The latter is externally provided by the IMF in our empirical setup.
@@ 3.1 Data and Forecast Targets
The WEO’s merit relative to other forecast sources has been established in studies such as Timmermann (2007) and Qu et al. (2024), which is why we consider it a promising candidate for this analysis. Started in 1990, the WEO has issued bi-annual forecasts for multiple (up to six) years into the future, for several variables including real GDP growth and inflation (International Monetary Fund, 2024a). We focus on the latter two variables throughout this paper. The code that downloads and processes the current WEO dataset from the IMF website is available through our GitHub repository (Becker et al., 2023).
@@
We focus on the 80% and 50% prediction intervals. As argued by Raftery (2016), an 80% interval satisfies the intuitive notion that a prediction interval should ‘typically’ cover the realizing outcome, while often being reasonably short (i.e., sharp) in practice. The $50\%$ interval is additionally constructed to provide further characterization of the forecast distribution for interested users, and to serve as a further check of the validity of the methodology.
@@ 3.2 Choice of Tuning Parameters
Table 5 and Figure 8 in the supplement summarize the training sample performance for the different combinations of the tuning parameters. Based on the performance in the training sample, and the considerations in Section 2, we use a rolling window of $R=11$ observations, the absolute error method, and the default empirical quantile method (type = 7) [^1] Note that for $R=11$ and the absolute error method, this is also equivalent to the inverse of the empirical distribution function, that is, type = 1. As we do not estimate quantiles at the tails of the distribution the arguments of Hyndman and Fan (1996) in favor of the type = 8 method do not apply for our setting. of the quantile function in R. Our training sample results indicate that shorter window lengths $R<11$ perform worse in terms of the IS, see Figure 8 in the appendix. The difference in performance is particularly pronounced for very short window lengths $74\leq R<$ 7 and at longer forecast horizons. By contrast, the benefit of increasing $R$ from $8$ to $11$ is remarkably small in all empirical setups (absolute and directional errors, all forecast horizons). Hence the marginal benefit of increasing the sample size seems to vanish rapidly as $R$ increases. We exclude choices $R>11$ since we cannot evaluate their performance given the time span of our training data, while an extension of the training data would have reduced the time span of our holdout sample, thus limiting our ability to assess out-of-sample performance. From a statistical perspective, using a short window length $R$ ensures adaptability to potential structural breaks (c.f. Inoue et al., 2017). We find that coverage rates on the training data are generally more favorable for the absolute than for the directional error method, sometimes on the order of about 10 percentage points (see Table 5 in the supplement), while interval scores are similar. We additionally present illustrative insights from the training period in Section B.3.1 in the supplement. Albeit attractive in principle, we argue that directional forecast errors sometimes display undesirable practical properties in our setup, and are thus not a good candidate to account for a potential skew in the distribution of forecast errors.
@@
Figure 2: Illustration of the process that we use to construct prediction intervals at the $50\%$ and $80\%$ confidence levels, using an existing point forecast and its history of forecast errors.
@@ 3.3 Benchmark Models
In order to put the performance of our proposed method in perspective, we compare it to several benchmarks. For ease of presentation, we focus on three main benchmark models in our main empirical analysis that each represent one of the following broader model types: a) univariate or b) multivariate time series models yielding different point predictions, where prediction intervals are calculated either based on past performance forecast errors as in our approach or based on c) the respective time series model specification. For further variants of these benchmark models and respective comparisons we refer to Section 3.5.3 below as well as Section B of the supplement. In particular, Tables 6 and 7 of the supplement provide comprehensive results in terms of evaluation metrics.
All considered benchmarks are of (vector) autoregressive time series form, which is one of the most common choices in macroeconomic forecasting. For example, Faust and Wright (2009) show that simple univariate autoregressions provide good forecasts of U.S. GDP growth when equipped with an appropriate information set, and Carriero et al. (2024) argue that vector autoregressive models with stochastic volatility compare favorably to nonlinear quantile regression methods for U.S. GDP growth, inflation and unemployment. We fit the main benchmark models to quarterly data from the OECD for inflation and GDP growth (see Section B.1 in the supplement), using expanding windows of training data. In order to roughly match the timing of IMF forecasts, we equip the model-based spring forecasts with data up until the first quarter, and the model-based fall forecasts with data up until the third quarter. As we argue in Section 3.5.1, evaluation with respect to IMF data is still justified.
@@
For the first main benchmark (denoted AR), we compute point predictions from an autoregressive model of order one, estimated via ordinary least squares. We construct a point forecast of the annual series as a weighted sum of quarterly forecasts and calculate prediction intervals from empirical quantiles of the model’s past forecast errors, analogous to our proposed procedure for the IMF forecasts. In Section 3.5.3 we also allow for a more general AR($p$) variant with data-driven choice of the lag length parameter $p$, and consider autoregressive models fitted to scarce annual IMF data.
For the second main benchmark (denoted BVAR), we consider a bivariate Bayesian vector autoregressive specification for inflation and GDP growth. As for the IMF forecasts (and for the AR model), we construct prediction intervals from empirical quantiles of the BVAR’s past forecast errors. Our BVAR implementation is based on the R package bvarsv (Krüger, 2015). While the latter considers the flexible model by Primiceri (2005) and Del Negro and Primiceri (2015), we restrict the priors in our main benchmark to shut off time variation in the model’s mean and covariance parameters, as in Knüppel et al. (2023, Section 5). We find that this simpler specification yields relatively stable parameter estimates for all countries and time periods. By contrast, Primiceri’s original priors that allow for time-varying parameters and stochastic volatility run into problems for some countries such as France and Great Britain after the extreme GDP growth rates observed at the onset of the Covid-19 pandemic. [^2] This problem could possibly be resolved by adjusting the model’s prior parameters to allow for an intermediate degree of flexibility, or by using a stochastic volatility specification that explicitly addresses Covid-19 outliers (as in Carriero et al., 2024). However, the details of such an adjustment would create many additional degrees of freedom in our setup, given that the severity of the Covid-19 outliers differs across the G7 countries. We hence use a simpler specification with constant parameters for our main analysis. Within the additional analyses discussed in Section 3.5.3, we also consider an approach that uses a stochastic volatility specification until 2019 and a constant volatility specification afterwards. This mixed variant entails a manual choice of regime switch. Even then, however, the performance of the model with constant parameters is largely similar with the exception of inflation at the next-year horizons where the mixed BVAR can outperform all methods and in particular our simple IMF-based approach.
@@
For the third main benchmark (denoted BVAR-direct), we use the same constant parameter specification as in BVAR, but compute prediction intervals directly from the model’s forecast distribution. That is, the forecast distribution is based on the model’s parametric assumptions, rather than on the empirical distribution of the model’s past point forecast errors. As is typical of Bayesian approaches, the BVAR-direct forecast distribution incorporates the effect of parameter estimation uncertainty. For completeness, Section 3.5.3 also considers “direct” variants of the other benchmark specifications.
Besides variations of the main benchmark models, Section 3.5.3 also contains further benchmark approaches such as models that use additional regressors and a simple ensemble.